OpenStudy (anonymous):

2x-5/2= 3x-1/4 How do you solve for x? and thank you in advance!

OpenStudy (anonymous):

Get all X terms on one side of the equation, and the numerical terms on the other: 2x -3x = 5/2 - 1/4 Add the terms: -x = 10/4 - 1/4 -x = 9/4 x = -9/4

OpenStudy (anonymous):

Thank you:)